### v1

# Resource-constrained Fairness

Abstract
Access to resources strongly constrains the decisions we make. While we might wish to offer every student a scholarship, or schedule every patient for follow-up meetings with a specialist, limited resources mean that this is not possible. Existing tools for fair machine learning ignore these key constraints, with the majority of methods disregarding any finite resource limitations under which decisions are made. Our research introduces the concept of “resource-constrained fairness" and quantifies the cost of fairness within this framework. We demonstrate that the level of available resources significantly influences this cost, a factor that has been overlooked in previous evaluations.
Keywords Fairness  $\cdot$ Constrained resources  $\cdot$ Responsible AI
1 Introduction
Machine learning models are used to make decisions in high-impact areas of our lives such as finance, justice, and healthcare (Mehrabi et al., 2021). Fair machine learning has emerged in response to the notion that simply making maximally accurate decisions is not enough and that training high-performance classifiers can result in both the transfer of existing biases from data to new decisions, as well as the introduction of new biases (Wachter et al., 2020).
Many studies that focus on improving fairness in machine learning overlook the practical limitations under which these models operate. For example, scenarios including university admissions, healthcare provision, and corporate hiring, are normally constrained by finite resources. Universities have a restricted quota of students to admit annually, healthcare facilities are bounded by available space and staff, and companies have a limited number of positions to fill. Even for banking, where in principle banks can keep making loans providing enough people pay them back, for any particular time period they will have a limited set of resources and only be able to loan out a certain amount. In all these cases, once the available resources are fully used, increasing the selection rate of disadvantaged groups must necessarily involve reducing selection from more privileged groups.
However, limited resources are not taken into account in current discussions in fair machine learning. Goethals et al. (2024) show that the number of selected individuals can vary substantially after deploying different bias mitigation methods, including learning fair representations (Zemel et al., 2013) and adversarial debiasing (Zhang et al., 2018). A couple of notable exceptions stand out: Kwegyir-Aggrey et al. (2023) discussed how practitioners need to adapt the threshold to ensure the outcomes meet their domain-specific needs, while Corbett-Davies and Goel (2018) explored constraining the number of selected instances to satisfy a certain budget.
In the standard setting of an unconstrained budget, Mittelstadt et al. (2023) posit that diminishing the performance of privileged groups solely to achieve fairness is not an optimal approach and refer to this as ‘leveling-down’. We agree with this viewpoint but argue that within contexts constrained by limited resources, some rebalancing is necessary to redistribute the resources. In these circumstances, deliberately not fully utilizing resources to ensure equality could also be termed ‘leveling down’.
In this study, we make the following contributions:
  • We reformulate the problem of fair machine learning as a resource-constrained problem, where we treat the positive decisions as resources to allocate among different groups.
  • We provide a theoretical connection to leveling up, demonstrating that under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
  • We quantify the cost of fairness in this resource-constrained framework. Our formulation allows us to examine the factors that affect this cost of fairness. We study the impact of factors inherent to the dataset, including overall uncertainty and per group base rate disparity, uncertainty, and size, alongside context-related factors such as the resource level and enforced fairness metric. We find that the selected positive decision rate has a substantial impact on this cost, an aspect that is not taken into account in previous evaluations. This failure to account for real-world constraints on fairness may be a contributing factor to the field’s lack of prior cases of enforcing fairness on production models, which contrasts starkly with the abundance of studies using fairness metrics solely for model testing and debugging.
  • We explore the practical implications of our findings and discuss what strategic actions practitioners might undertake to mitigate trade-offs between groups when fairness is constrained by resource limitations.
2 Background
We consider fairness in classification. Starting with a classifier $c_{w}(\cdot)$ parameterized by weights $w$. Let $\mathrm{H}_{\cal D}[c_{w}]$be a measure of expected harm of a classifier $c_{w}(\cdot)$ over a particular distribution $\cal D$ where $y_{x}$ represents the true target label of instance $x$, i.e.,
$\mathrm{H}_{\cal D}[c_{w}]=\mathrm{E}_{x\in\cal D}H(y_{x},c_{w}(x))$ (1)
Such measures of a harm $H$ might be $1-\text{precision}$ (for example, when measuring the proportion of people incorrectly stopped by the police); or $1-\text{recall}$ (when measuring the number of cancer cases that go unflagged for follow-up treatment).
When we are concerned about fairness in classification, we usually want to measure fairness with respect to a protected attribute, such as gender or ethnicity. Using this protected attribute, we can partition the dataset into groups. Group fairness metrics measure (in)equality between these groups with respect to some measure of harm. For example, equal opportunity requires the recall between groups to be equal, while demographic parity enforces an equal selection rate (Verma and Rubin, 2018). Typically, a fair classifier is found by minimizing some global loss $\ell$ (such as accuracy or a continuous proxy such as the logistic-loss) while ensuring that the harm is the same per group.
This means that we are searching for a solution to the following problem (where $\cal G$ is a partitioning of the distribution into groups with respect to a particular protected attribute):
$\min_{w}\,\mathrm{E}_{x\in\cal D}\ell(y_{x},c_{w}(x))\text{ such that }$ (2)
$\mathrm{H}_{g_{1}}[c_{w}]=\mathrm{H}_{g_{2}}[c_{w}]\,\forall g_{1},g_{2}\in\cal G$
2.1 Leveling down
Mittelstadt et al. (2023) observe that methods to enforce fairness often level-down; that is, they may enforce fairness by decreasing harm in some groups, but also by increasing harm to other groups (e.g. naively enforcing Equal Opportunity while detecting cancer will often result in some groups receiving a lower rate of cancer detection than they would otherwise). Note that this process of enforcing fairness can alter the overall selection rate of the model (Goethals et al., 2024), and is distinct from any leveling down that might be inherently required by the resource constraints discussed earlier. To this end, they suggested replacing Equation LABEL:eq:fair_class with rate constraints, which enforced that the harm in question (e.g. being denied follow-up care) should be below a certain level $k$ for every group:
$\min_{w}\,\mathrm{E}_{x\in\cal D}\ell(y_{x},c_{w}(x))\text{ such that }\mathrm{H}_{g}[c_{w}]\leq k\,\forall g\in\cal G$ (3)
This is the same as saying that in the worst case, the harm should be below k, for any group. The challenge now becomes, “How should $k$ be set?” When deciding to deploy a particular model, stakeholders and data scientists often have to agree on acceptable global levels of harm, for example an acceptable recall rate (referred to as sensitivity in the medical literature) for early cancer detection. A similar process can be used to select $k$ per group.
This notion of leveling-up, and decreasing the harm in Equation 3 is related to but distinct from minimax-fair machine learning, which minimizes the error for the worst-off group (Martinez et al., 2020). In both minimax fairness and leveling-up the concern is with reducing the harm to the worst-off group and only increasing harms to other groups when necessary, but in minimax fairness the harm is exclusively assumed to be caused by a lack of accuracy or high log-loss (this leads to a formulation of minimize the maximal loss, hence minimax). In comparison, leveling-up is concerned with decreasing other harms such as per-group recall or selection rate, which are distinct from accuracy.  [^1] Both notions are closely aligned to the philosophical principle of the ‘maximin rule’ (Rawls, 2017). According to Rawls, resources should be distributed such that they maximize the benefits for the least advantaged members of society. However, as a fundamental limitation, neither minimax nor leveling-up considers what should be done under constrained resources.
Taking a step back, we can ask the questions: “If $H$ is a harm, why not set $k$ to zero? If particular groups are being harmed by not being selected, why not select everyone, and not bother with machine learning? Similarly, if people are being harmed by a failure to detect cancer, why not schedule everyone for follow-up testing? ”
There are multiple possible answers here. One is a matter of personal utility, that harms are generally not one-sided. Failure to repay a loan can lead to bankruptcy for both the lender and customer, resulting in devastating personal consequences. Scheduling unnecessary medical tests is at best alarming, and in the worst case can result in death, depending on how intrusive the follow-up tests are. These types of considerations have been addressed by prior work on the cost of fairness, where personal utility is measured by accuracy, (Menon and Williamson, 2018) and fairness on the ground, which measures the effect of the actual implementation of fairness principles in real-world situations (Bakalar et al., 2021). Ideally, we would prefer to offer scholarships to every student and fast-track treatment for every patient. However, real-world scenarios often lack the necessary resources to achieve this. Understanding how fairness choices are limited and guided by resource constraints is thus crucial to move fair machine learning from a testing to production environment.
Fair classification under limited resources is also related to the domain of fair allocation within welfare economics. This field focuses on ensuring that resources such as time or physical goods are distributed amongst actors in a way that meets certain criteria. This leads to a tension between maximum utility and fairness (Bertsimas et al., 2011, 2012; Donahue and Kleinberg, 2020). Elzayn et al. (2019) describe a resource allocation problem as a problem where there is a limited supply of resources to be distributed across multiple groups with different needs, such as allocating loans or disaster response. We suggest that when the decision-maker has limited resources, all fair classification tasks can be interpreted as resource allocation problems, where positive decisions represent the resources to be distributed.
2.2 Constrained resources
We formalize the notion of resources, as the number of instances predicted as positive by the machine learning model. This term can be used interchangeably with capacity, or with selection rate or positive decision rate when speak about the proportion of positively predicted instances. We define $\mathrm{R}_{\cal D}[c_{w}]$to be the selection rate (i.e. the expected proportion of positive decisions) of a classifier over a distribution $\cal D$.
Proposition 1.
Under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
Proof.
We consider the problem of optimizing the distribution of limited resources to minimize the maximum harm experienced by any group, given by:
$\min\max_{g}\mathrm{H}_{g}[c_{w}]\text{ such that }\mathrm{R}_{\cal D}[c_{w}]\leq R$ (4)
where $\mathrm{R}_{\cal D}[c_{w}]$is a weighted sum of group-specific rates $\mathrm{R}_{g}[c_{w}]$ with all weights positive, thus making $\mathrm{R}_{\cal D}[c_{w}]$strictly increasing with respect to $\mathrm{R}_{g}[c_{w}]$.
We first consider the case where $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing as a function of $\mathrm{R}_{g}[c_{w}]$ (for example, if the harm is $1-\text{recall}$, we expect the harm to be strictly decreasing when more individuals of the group are selected). As $\mathrm{H}_{g}[c_{w}]$ can be written as an invertible function of $\mathrm{R}_{g}[c_{w}]$, $\mathrm{R}_{g}[c_{w}]$ is also strictly decreasing with respect to $\mathrm{H}_{g}[c_{w}]$. Since $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with $\mathrm{R}_{g}[c_{w}]$, and $r_{D}$ is strictly increasing with respect to $\mathrm{R}_{g}[c_{w}]$, it follows that $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with respect to $r_{D}$.
Define $W$ as the set of worst-off groups under the harm $H$, such that:
$W=\arg\max_{g}\mathrm{H}_{g}[c_{w}]$ (5)
The optimum must occur when $\mathrm{R}_{\cal D}[c_{w}]=R$. If not, one could distribute the remaining resources such that the rates $\mathrm{R}_{g}[c_{w}]$ for the worst-off groups increase without exceeding R, thus reducing the maximum harm (since $\mathrm{H}_{g}[c_{w}]$ is decreasing with respect to $\mathrm{R}_{g}[c_{w}]$).
Furthermore, the optimum is reached when the harm levels are as equal as possible across all groups (so all groups are in $W$). If this were not the case, then there would be at least one group not in $W$, that could sacrifice some of its resources to the worse-off groups. Adjusting the allocation in such a way that the group not in $W$ receives less, while groups at maximal harm receive more, would lead to a reduced overall maximum harm due to the strict decrease of $\mathrm{H}_{g}[c_{w}]$ with $\mathrm{R}_{g}[c_{w}]$. Thus, under these conditions, equality in harm distribution is enforced in the optimal solution, ensuring that the distribution of resources maximally benefits the least advantaged groups within the constraints set by $\mathrm{R}_{\cal D}[c_{w}]\leq R$.
This proof also holds when the harm considered is strictly increasing with respect to $\mathrm{R}_{g}[c_{w}]$ (e.g., a harm corresponding to $1-\text{precision}$). In this case, the optimal solution is trivial with an overall selection rate of 0, and equality will be preserved. However, in this case, it might be more realistic to flip the sign and require the overall selection rate to be at least some R. We can simply replace the classifier $c_{w}$ with its negation, and now find that the harm is strictly decreasing with respect to the selection rate of the new classifier. By revisiting the previous proof, we find that equality must hold.
Hence, we find that equality must hold in all four cases (strictly increasing or strictly decreasing harm, and a global selection rate that is either constrained above or below). ∎
2.3 Cost of fairness
Before we proceed further, we should note that assigning a cost to actions is inherently a political action, and that often these costs reflect the beliefs of data scientists and other professionals as much as they do the raw data. This is particularly apparent in the case of personal loans.
While banks have good knowledge of the repayment habits of people they have lent money to in the past, they have much less knowledge of how people they did not lend money to might repay. In no small part, this knowledge asymmetry has allowed for the persistence of institutional redlining whereby loans are routinely denied to people in black or Hispanic dominant neighbourhoods when they would be granted to individuals in similar circumstances that live in white dominant neighbourhoods  (Markup, 2021). Similarly, while it is tempting to assume that medical data is reliable, and for example, that we are forecasting the result of a future biopsy that will infallibly determine if cancer is present or not, for many problems the information is less clear-cut, and ground-truth may come from two experts attempting to diagnose from limited post-mortem records, with a third expert adjudicating in the case of disagreement Lee et al. (2023).
Indeed, one justification for constraints such as demographic parity is that ground-truth data is often gender- or ethnically-biased, and in some circumstances we can a priori expect uniform rates across these populations (in line with the “We’re All Equal” world-view of Friedler et al. (2021), which asserts that there are no innate differences between groups). As such, any estimate of cost comes with the usual caveat of assuming the ground-truth is correct. But even as simplified approximations, these costs remain useful for understanding the potential trade-offs in enforcing fairness.
With this in mind, we consider the following question:
What is the global change in harm from an optimal classifier when we minimise the harm of the worst-off group?
We have shown that under constrained resources, this leads to the same results as enforcing fairness, i.e. equality of harms between groups. We can measure this by the difference between the harm of the optimal classifier $c^{o}_{w}(.)$ and the classifier that satisfies fairness $c^{f}_{w}(.)$.
$\mathrm{H}_{\cal D}[c^{f}_{w}]-\mathrm{H}_{\cal D}[c^{o}_{w}]$
While other studies have previously analyzed the cost of fairness (Friedler et al., 2019; Haas, 2019; von Zahn et al., 2021), they did not consider the implicit trade-off that comes from constrained resources in decision-making, but instead measured how classifiers deteriorate with fairness.
In some cases, we might also be interested in other performance metrics than the fairness metric we enforce. If we want to quantify the loss in precision, we could ask:
What is the change in precision if we minimize the harm of the worst-off group?
This represents a common scenario. In a medical context it corresponds to the question what proportion of healthy patients instead of sick will we see if we increase test sensitivity for the worst off groups? We measure this cost and the factors that contribute to it in our results.
Both questions can also lead to a follow-up question:
What is the increase or decrease in selection rate needed to preserve the current rate of global harm, if we minimise the harm of the worst off-group?
The last question represents a common political solution to this problem. When particular groups are disadvantaged by the status quo, it is often easier to increases the resources available, and target these resources at the disadvantaged groups, rather than requiring currently advantaged groups to accept less access to resources. The number of available resources can be an alterable parameter that should be taken into account.
3 Materials and methods
3.1 Materials
We use several real world datasets that are common in the domain of fair machine learning (Le Quy et al., 2022). Besides these often used datasets from the tabular domain, we also extend our findings to two datasets (CelebA Liu et al. (2015) and Fitzpatrick17k Groh et al. (2021)) from computer vision. The Adult Income dataset comprises information from the 1994 census data, focusing on whether an individual’s annual income surpasses \$50,000. The Compas dataset gathers demographic details and criminal records of defendants from Broward County to forecast the likelihood of reoffending within two years. The Dutch Census dataset from 2001 captures aggregated demographic data in the Netherlands, utilized to determine if an individual’s occupation falls into a high-level (prestigious) or low-level category. The Law Admission dataset contains data from a 1991 survey by the Law School Admission Council (LSAC) across 163 U.S. law schools, aimed at predicting a student’s success on the bar exam. The CelebA dataset comprises over 200,000 celebrity images, each annotated with 40 attribute labels ranging from hair color to emotions, along with 5 landmark locations. We use the attribute ‘wearing earrings’ as target label, as it is highly skewed towards the non Male class Wang et al. (2020) and thus has a significant base rate disparity. The Fitzpatrick17k dataset consists of approximately 17,000 dermatologist-curated skin lesion images, categorized across the Fitzpatrick skin type scale. We prepossess the data and binarize the labels following the approach of Zong et al. (2022). We represent each dataset and specify the protected and target attributes Table 1. We also report the base rate disparity for each dataset, which is defined as the actual difference in the proportion of positive outcomes between groups in a dataset. If we consider a binary sensitive attribute $S$, with $s$ representing the protected group and $ns$ the privileged group, then it can be mathematically expressed as:
$P(Y=1\mid S=ns)-P(Y=1\mid S=s)$
Table 1: Used datasets. The values between brackets represent the percentage of the dataset with that value.
[…]
3.2 Methods
3.2.1 Machine learning model
For the tabular datasets, we use the eXtreme Gradient Boosting (XGBoost) model, a gradient boosting framework optimized for both speed and performance Chen and Guestrin (2016). We optimize the number of boosting rounds through 5-fold cross-validation on the training set with early stopping after 10 rounds if no improvement is seen. The optimized model is trained on the entire training set and evaluates the probability of positive class outcomes. For the image datasets, we employ a Resnet-50 (CelebA) and Resnet-18 (Fitzpatrick-17k) backbone He et al. (2016) pretrained on ImageNet Deng et al. (2009) for feature extraction.
3.3 Metrics
3.3.1 Performance metrics
We want to evaluate the performance of the default allocation and the fair allocation over a range of selection rates (where the number of selected instances is the same for both allocations). When the selection rate of a machine learning model is fixed, it implies that a set proportion of the instances will be chosen based on the highest scores predicted by the model, regardless of their actual scores. A useful performance metric to measure is the proportion of relevant instances among the top $R$ instances selected by the model (where $R$ denotes the resource level). This means the number of selected instances that actually have a positive target label.
Precision at R=Number of actual positives in the top $\text{Precision at $R$}=\frac{\text{Number of actual positives in the top $R$}}{R}$
While precision at the top $R$ relates to the efficient use of limited resources, knowing this value also determines the values of both recall at $R$ and accuracy at $R$ for a specific dataset and a fixed value of $R$ (Rodolfa et al., 2021). We measure the cost of fairness as the loss in precision when using the fair allocation instead of the optimal allocation. For the sake of completeness, we also report the cost as loss in recall (which measures the proportion of actual positives in the top $R$ over the total number of actual positives) in Table 2 and Figure 6.
Besides the binary predictor $\hat{Y}$, each classifier will also output a prediction score S, with the interpretation that higher values of S corresponds to a greater likelihood of Y = 1 (Hardt et al., 2016). This score is transformed in a binary prediction by using a threshold: $\hat{Y}=\mathds{1}\{S>t\}$. As we use the classifiers at various thresholds, it makes more sense to evaluate the overall performance of the classifier by using the prediction scores, as this is unaffected by the choice of threshold. This is measured by the Area Under the ROC Curve (AUC). The formula for the AUC score:
$P(S(x_{i})>S(x_{z})|y_{i}=1,y_{z}=0)$
This formula measures the probability that the classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance.
3.3.2 Fairness metrics
We focus on two widely recognized fairness measures used to assess disparities between groups.
Demographic parity (also known as statistical parity) requires that the rate of positive decisions is roughly equal for both the protected group and the privileged group:
$P(\hat{Y}=1|S=s)\approx P(\hat{Y}=1|S=ns)$
Equal opportunity requires the true positive rate to be approximately the same across groups (Hardt et al., 2016), which enforces equal recall:
$P(\hat{Y}=1|S=s,Y=1)\approx P(\hat{Y}=1|S=ns,Y=1)$
3.4 Experimental set-up
Figure 1: Experimental framework
Figure 1 illustrates our experimental framework. We use a standard train-test split, where the model is trained on the training set and the predictive performance and cost of fairness is evaluated on a separate test set.
As discussed, machine learning models do not only output prediction labels but also prediction scores. We use these scores to generate the prediction labels for a range of selection rates. For the ‘default’ labels, we select the top $R$ of instances based on the highest prediction scores, where $R$ is the available resource level, and $r$ the resulting selection rate. For the ‘fair’ labels, we compute the proportion of resources to be allocated to each group. Allocating resources for demographic parity is relatively straightforward, as we select the same percentage in each group (illustrated in Figure 2). For equal opportunity, it involves calculating the threshold needed to equalize the number of true positives. [^2] Note that the thresholds to enforce demographic parity can be calculated on the test set, but the thresholds for equal opportunity should be calculated on a separate validation set, as they require access to the target label. To ensure that the number of instances predicted as positive is equal to the available resource level, we do not calculate the thresholds but the proportions. We calculate the proportion of resources to allocate to each group on the validation set, and then enforce this on the test set for the resource level $R$. This set-up belongs to the class of post-hoc bias mitigation strategies, that attempt to make the output of machine learning model fair after the model has been trained (Mehrabi et al., 2021). We conduct this analysis across a complete range of selection rates $r$ (from 1% to 100 %) to represent different resource levels $R$.
Figure 2: Example of implementing demographic parity in a resource-constrained set-up. If we want to satisfy demographic parity for a given resource level R, we need to switch some of the selected instances from the privileged group for instances of the protected group.
The main objective of our work is to quantify the “cost of fairness", defined as the difference in performance (measured as either precision or recall) between the default and fair labeling approaches, for a fixed number of selected instances ($R$). We vary the level of available resources R, to investigate how this influences the cost of fairness.
4 Results
4.1 What influences the cost of fairness?
Table 2: Model performance (entire population, privileged group and protected group) and the average cost of fairness (measured by the loss in precision and recall) when enforcing demographic parity and equal opportunity on all datasets. The average is calculated over all the selection rates from $1\%$ to $100\%$.
Dataset Adult Compas Dutch Law CelebA Fitzpatrick17K
AUC 0.899 0.811 0.917 0.871 0.957 0.826
$AUC_{priv}$ 0.884 0.797 0.884 0.852 0.931 0.829
$AUC_{prot}$ 0.887 0.793 0.914 0.842 0.969 0.814
Avg. loss in precision (DP) 0.024 0.025 0.028 0.006 0.069 0.001
Avg. loss in precision (EO) 0.010 0.013 0.007 0.004 0.009 0.000
Avg. loss in recall (DP) 0.024 0.016 0.028 0.004 0.062 0.001
Avg. loss in recall (EO) 0.012 0.007 0.008 0.003 0.013 0.001
In Table 2, we analyze the AUC performance of our machine learning model for different groups: the entire population, and separately for protected and privileged groups. Additionally, we report the average cost of enforcing fairness as the loss in both precision and recall between the default allocation and the fair allocation, averaged over all selection rates.
We see that the average cost of enforcing fairness is the lowest for the Fitzpatrick17K dataset, which also has a very low base rate disparity. The average cost is also very low for the Law dataset, despite its significant base rate disparities. A possible reason for this could be the relatively small size of the protected group in this dataset, which results in less reallocation of resources. The average cost of enforcing fairness across other datasets is more similar, however it is worth noting that the cost associated with enforcing both fairness metrics on CelebA is higher than for the other datasets. We see that the average loss in precision and in recall of all datasets is in line with each other, although the exact numbers differ.
It is hard to attribute the difference in costs between the datasets to a single factor, as many of the parameters will be different. This is why we perform a separate analysis on the Adult Income dataset. Here, we systematically vary key parameters—one at a time—to observe their effects on the cost of fairness. Specifically, we explore modifications to the base rate disparity, the size of the protected group, the noisiness of the whole dataset and the noisiness of the protected group. The impacts of these changes are visualized in Figure 3, with the original dataset’s results represented by a black line in each figure.
(a) Base rate disparity (DP)
(b) Base rate disparity (EO)
(c) Noise (DP)
(d) Noise (EO)
(e) Subgroup noise (DP)
(f) Subgroup noise (EO)
(g) Subgroup size (DP)
(h) Subgroup size (EO)
Figure 3: Impact of parameters on the cost of fairness, measured as the loss in precision (Adult Income dataset). We see that reducing the base rate disparity leads to a lower cost of fairness. Similarly, introducing more noise across all groups, also lowers the cost of fairness, but introducing more noise exclusively in the protected group, increases the cost of fairness. Finally, reducing the size of the protected group, results in a decrease in the cost of fairness as well.
(a) Adult
(b) Compas
(c) Dutch
(d) Law
(e) CelebA - Earrings
(f) Fitzpatrick17k
Figure 4: Cost of fairness (loss in precision) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). We see that the cost of fairness for both metrics heavily depends on available resource level and thus the used selection rate.
4.1.1 Base rate disparity
First, we alter the base rate disparity. We see in Figures 3(a) and 3(b), that a reduction in the base rate disparity leads to a lower cost of fairness. This trend is more pronounced when enforcing DP compared to EO, but it is evident under both fairness metrics. Menon and Williamson (2018) also find that the trade-off between accuracy and fairness is determined by the strength of the correlation between the sensitive attribute and the target variable.
4.1.2 Noise
We alter the performance of the model by introducing random noise to the feature values of the whole dataset. The base rates of both groups remain the same, but the model will have more trouble with correctly classifying all individuals, which results in a lower AUC score. We add noise with varying degrees from 0 to 0.5 as depicted in Figures 3(c) and 3(d). We see that when the model performance goes down, the average cost of fairness also goes down, both for DP and EO. This decrease shrinks the gap between the performances of each group, and hence reduces the overall cost of fairness. This phenomenon might also explain why the average cost of fairness is so high for the CelebA dataset in Table 2. The model is very good in distinguishing negatives and positives from each other (as measured by the AUC), so the average cost will be high. [^3] Following this reasoning, we see that if we would fit a perfect model (so the model predicts the target label perfectly for all instances) to a dataset with some level of base rate disparity, then the maximum cost would be reached at a point between the base rates and be exactly equal to the level of base rate disparity in that dataset. In the case of CelebA, the model is not perfect but the maximum cost is also relatively close to the level of the base rate disparity.
4.1.3 Subgroup noise
We can also add noise only to the members of the protected group. This will lower the performance of the model for individuals of the protected group, but not for individuals of the privileged group. As shown in Figures 3(e) and 3(f), this modification increases the cost of fairness for both DP and EO, with a larger effect for EO. This effect is also larger at lower selection rates. Chen et al. (2018) also find that when there is a difference in noise level, and available covariates are not equally predictive of the outcome in both groups, fairness cannot be satisfied without sacrificing accuracy. Dutta et al. (2020) confirm that if there is not enough separability information for one group compared to the other, being fair will reduce the accuracy.
4.1.4 Size of the protected group
Lastly, the size of the protected group can significantly influence the cost of fairness. When the protected group is smaller, fewer adjustments are necessary to achieve fairness, thereby reducing the cost. This relationship is clearly illustrated in Figures 3(g) and 3(h), where reducing the size of the protected group, achieved by randomly excluding part of this group, consistently lowers the cost of fairness. This is a logical finding, however, we did not see it being discussed in literature before.
4.2 What is the impact of the available level of resources (or selection rate)?
Figures 4(a)-4(f) demonstrate the loss in precision for various selection rates when using a fair allocation compared to the default allocation of the machine learning model. We see that this cost varies substantially depending on the level of available resources. For the Adult (Figure 4(a)), Compas (Figure 4(b)), CelebA (Figure 4(e)) and Fitzpatrick 17K (Figure 4(f)) datasets, the cost is highest for low selection rates, while for the Dutch (Figure 4(c)) dataset, the cost is the highest for medium selection rates, and for the Law (Figure 4(d)) dataset, the cost is the highest for high selection rates. The influence of resource level on the cost of fairness has not been discussed before, yet it is a critical factor. Depending on the resource level, the cost can either be negligible or very high. The loss in recall can be seen in Figure 6. We see that the patterns are in line with the loss in precision. We can also partially agree with Rodolfa et al. (2021) who posited that for very constrained top-$k$ settings, so when the selection rate is very low, the cost of fairness can be negligible. However, we only find this to be true when the base rate is a lot higher and the performance of the model is good enough. When the base rate is low, the cost of fairness will usually be high for low-resource settings.
(a) Selection rate = 10 %
(b) Selection rate = 25 %
(c) Selection rate = 50 %
(d) Selection rate = 75 %
Figure 5: Allocation results for different resource levels (Adult Income dataset). Left point on the axis represents the situation where the resources are maximally allocated to the privileged group, while the right point represents the situation where the resources are maximally allocated to the protected group.
We can get more insights into how these trade-offs behave for each resource level by analyzing the different scenarios. For the Adult dataset, this is shown in Figures 5(a)-5(d). These figures demonstrate how different allocations lead to different levels of precision for one fixed resource level. The two extreme points of the curve represent all the resources being awarded to either the privileged group (red dot on the left) or the protected group (purple dot on the right). However, also for these extreme points, when the total level of resources is higher than the size of one of the groups, some of the resources will still be awarded to the other group (as can for example be seen in Figure 5(c)). We see that the optimum is not reached by awarding all the resources to one of the groups, but somewhere in the middle (darkblue dot). We note that the precision of the unconstrained (‘unfair’) model, will be very close to the optimal allocation, and that the allocations required by the fairness metrics (both DP and EO) will lead to a lower precision for every resource level (Figures 5(a)-5(d)). However, this difference is a lot larger for low resource levels (Figures 5(a)-5(b)). In Figure 5(c), the default allocation of the machine learning model is very unfair, but using the fair allocation (both DP and EO) results in a very low difference in precision. This demonstrates that although the initial model allocation is very unfair, this does not necessarily result in a high cost of fairness. In Figure 5(d), the default allocation of the machine learning model is already approximately fair, and hence the cost of fairness is also low.
4.3 What is the impact of the chosen fairness metric?
We can see in Table 2, and in Figures 4 and 6 that the cost of fairness is usually lower when enforcing equal opportunity than when enforcing demographic parity. This is because the number of instances that needs to be ‘swapped’ is typically lower when enforcing equal opportunity. Liu et al. (2018) also find that the selection rate enforced by equal opportunity is likely to be much closer to the optimal than the selection rate enforced by demographic parity, while Hardt et al. (2016) show that enforcing equal opportunity leads to less loss in profits than enforcing demographic parity.
The setting where EO would results in a higher number of swapped instances, is when an equal true positive rate results in stricter requirements than an equal positive rate. A case where this could happen is when one of the groups has significantly lower data quality. Demographic parity suffers less under this scenario, as here we do not care if the predictions are accurate or not. However, to enforce equal opportunity, the algorithm must do equally bad on all groups, leading to a worse overall performance. Indeed, as shown in Figures 3(e) and 3(f), the cost of enforcing Equal Opportunity (EO) increases significantly more rapidly compared to the cost of enforcing Demographic Parity (DP) when one subgroup becomes noisier.
It is important to note that in our implementation, we enforce fairness perfectly. Many papers study a relaxed formulation of fairness constraints, where some level of unfairness is allowed (Liu et al., 2018; Friedler et al., 2019; Haas, 2019).
4.4 Summarizing the costs
We can approximate the overall costs as follows: the cost of fairness depends strongly on two factors (i): the size of the disadvantaged group (if they are small, treating them differently will have little effect on global measures). (ii) The performance gap between the subsets of data that are treated differently under fairness (i.e. those individuals in the protected group that gain an advantage because of the use of fairness, and those that become disadvantaged in the privileged group due to fairness). This second factor explains much of what we see. For example, the increase of global noise decreases performance for both groups, at broadly similar proportions. This overall decrease shrinks the gap between the per-group performance, reducing the overall cost of fairness. Similarly, increasing classification noise for the worst performing group or increasing the difference in base rates increases the cost of fairness.
Another key finding is that the cost is typically bell-shaped with respect to the global selection-rate. The highest costs occur when the selection rates are closest to the selection rates of an unconstrained classifier. This means that many of the existing works on the cost of fairness (Menon and Williamson, 2018; Friedler et al., 2019; Hort et al., 2023) will overestimate the cost of fairness for resource constrained scenarios. For scenarios where the selection rate substantially exceeds the base-rate, this is because for any informative classifier, the majority of candidates that might be positive would be selected early on, regardless of which group they belong to, and the remaining pool of candidates are likely to have negative labels regardless of which group they belong to. Similar behaviour occurs with low selection-rates. If a classifier is correctly confident that a small subset of each group do take a positive label then fairness can be enforced using these subsets for little cost. However, it must be emphasized that this behaviour is not mathematically guaranteed, and if the classifier response is genuinely uninformative within each group, then the cost of fairness will be constant for all selection rates that are below the proportion of individuals belonging to the advantaged group.
5 Discussion
So far, we measured the cost of fairness as the loss in precision by measuring the difference between the fair allocation and the default allocation of the machine learning model. However, this does not constitute the actual cost, as this is connected to the tangible impact of each correct or erroneous prediction. Consider the scenario where an algorithm can select 10,000 instances, with a loss in precision of 0.01 due to fairness constraints. This adjustments results in 100 fewer true positives and true negatives, consequently increasing the false positives and false negatives by the same number. In the context of healthcare screening, this translates to 100 missed cancer diagnoses, and 100 unnecessary follow-up tests.
Strategic adjustments may become necessary depending on the outcomes of this cost-benefit analysis. Up until now, we assumed the resource level to be fixed, but practitioners may find it feasible to increase or decrease this level, particularly if the resource levels were initially determined under an unfair model allocation. For instance, in healthcare, additional investments could be directed towards expanding screening facilities to counterbalance the decrease in true positives, effectively maintaining the detection rates prior to implementing fairness measures. Similarly, in educational settings such as student admission, institutions might adjust the size of admitted cohorts, either decreasing it to preserve the average quality, or increasing it to have the same number of graduates. Additionally, after doing this analysis, one could just select a resource level where the cost of fairness is low. Donahue and Kleinberg (2020) also discuss how increasing the level of available resources is a critical goal where advocacy and political action can play a key role. However, it will often not be possible to have a resource level equal to the whole population (Donahue and Kleinberg, 2020).
Another strategic approach could involve refining the precision for the protected groups by collecting more data or enhancing data quality, albeit at an initial cost. This investment might reduce the overall fairness cost, yielding a more profitable model over time. However, high quality data may simply not be available for all groups or infeasible to collect for a variety of legal, political, and practical reasons, meaning it cannot be assumed that collecting more or “better" data will necessarily solve inequalities in resource distribution (Perez, 2019; Pot et al., 2019; Wachter et al., 2020).
6 Conclusion and limitations
In this work, we argued that fairness in machine learning should be investigated as a resource-constrained problem. We introduced the notion of resource-constrained fairness, and quantify the cost of fairness for a fixed level of resources (or a fixed selection rate). This allows us to investigate the actual cost of fairness, and not changes in performance metrics that are heavily influenced by a change in the overall selection rate. Our analysis closely aligns with the real-world challenges in enforcing fairness, where resources are often fixed, and allows organizations to have reasonable expectations about what costs to expect and why. Furthermore, we demonstrate that several dataset properties are relevant for the cost of fairness and that the cost of implementing fairness measures significantly depends on the decision-making context, particularly the used fairness metric and the level of available resources. This highlights the importance of evaluating the cost of fairness within the specific context of your problem.
However, we also see various limitations to our work. First, as already mentioned, assigning a cost to enforcing fairness is essentially a value judgement. We calculate the cost under the assumption that the ground truth is correct, while in reality, these labels can be biased as well. Despite this assumption, our results remain useful to understand potential trade-offs in enforcing fairness. Lastly, in our set-up we assume access to a static sensitive attribute. In reality, access to these attributes might be hard to obtain, as legal and ethical constraints may impose constraints on obtaining or utilizing certain sensitive information (Haeri and Zweig, 2020; Veale and Binns, 2017). In these scenarios, inferred characteristics could be used during deployment (Lohaus et al., 2022). Despite these limitations, this work still makes important contributions towards the field of Responsible AI in practice, by taking into account real-world constraints.
Acknowledgments
This work has been supported through research funding provided by the Wellcome Trust (Grant number 223765/Z/21/Z), Sloan Foundation (Grant number G-2021-16779), Department of Health and Social Care, EPSRC (Grant number EP/Y019393/1), and Luminate Group. Their funding supports the Trustworthiness Auditing for AI project and Governance of Emerging Technologies research programme at the Oxford Internet Institute, University of Oxford. Sofie Goethals was funded by Flemish Research Foundation (Grant number 11N7723N), and by a travel grant (Grant number V413024N) awarded by Flemish Research Foundation to visit the University of Oxford.
Appendix
(a) Adult
(b) Compas
(c) Dutch
(d) Law
(e) CelebA - Earrings
(f) Fitzpatrick17k
Figure 6: Cost of fairness (loss in recall) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). We see that the cost of fairness for both metrics heavily depends on available resource level and thus the used selection rate.