### v4

# Resource-constrained Fairness

Abstract
Access to resources strongly constrains the decisions we make. While we might wish to offer every student a scholarship, or schedule every patient for follow-up meetings with a specialist, limited resources mean that this is not possible. When deploying machine learning systems, these resource constraints are simply enforced by varying the threshold of a classifier. However, these finite resource limitations are disregarded by most existing tools for fair machine learning, which do not allow the specification of resource limitations and do not remain fair when varying thresholds. This makes them ill-suited for real-world deployment. Our research introduces the concept of “resource-constrained fairness" and quantifies the cost of fairness within this framework. We demonstrate that the level of available resources significantly influences this cost, a factor overlooked in previous evaluations.
Keywords Algorithmic Fairness  $\cdot$ Responsible AI in practice
1 Introduction
Machine learning models are used to make decisions in high-impact areas of our lives such as finance, justice, and healthcare (Mehrabi et al., 2021). Fair machine learning has emerged in response to the notion that simply making maximally accurate decisions is not enough and that training high-performance classifiers can result in both the transfer of existing biases from data to new decisions, as well as the introduction of new biases (Wachter et al., 2020).
Many studies that focus on improving fairness in machine learning overlook the practical limitations under which these models operate. For example, scenarios including university admissions, healthcare provision, and corporate hiring, are normally constrained by finite resources. Universities have a restricted quota of students to admit annually, healthcare facilities are bounded by available space and staff, and companies have a limited number of positions to fill. Even for banking, where in principle banks can keep making loans providing enough people pay them back, for any particular time period they will have a limited set of resources and only be able to loan out a certain amount. In all these cases, once the available resources are fully used, increasing the selection rate of disadvantaged groups must necessarily involve reducing selection from more privileged groups.
Figure 1: Illustration of how the different bias mitigation methods enforce fairness at a specific selection rate (or threshold) but break when we consider the full range of possible selection rates (Adult Income dataset) . [^2] The icons represent the labels of the bias mitigation methods at one specific threshold. The dashed line represent the output for the other thresholds, by using the prediction scores generated by that method. Note that this approach is not possible for Fairlearn as they do not return a prediction score.
However, limited resources are not taken into account in most discussions in fair machine learning. Kwegyir-Aggrey et al. (2023) also highlight how practitioners typically need to adapt the threshold to ensure that outcomes meet their domain-specific needs, while corbett2023measure argue that the trade-offs in fair decision-making are most acute when there are constrained resources. This failure to account for real-world constraints on fairness may be a contributing factor to the field’s lack of prior cases of enforcing fairness on production models, which contrasts starkly with the abundance of studies using fairness metrics solely for model testing and debugging.
Figure 2 illustrates how demographic parity, one of the most common notions of fairness, varies with selection rate. Many of the currently popular bias mitigation methods, enforced using a range of tools from AIF360 Bellamy et al. (2018) and Fairlearn Bird et al. (2020), aim to ensure fairness at a specific threshold. However, as the thresholds vary, fairness oscillates wildly, revealing that these methods may break when considering the whole range of possible selection rates.
In the standard setting of an unconstrained budget, Mittelstadt et al. (2023) posit that increasing harm to privileged groups solely to achieve fairness is not an optimal approach and refer to this as ‘leveling down’. We agree with this viewpoint but argue that within contexts constrained by limited resources, some rebalancing is necessary to redistribute the resources. In these circumstances, deliberately not fully utilizing resources to ensure equality could also be termed ‘leveling down’. .
In this study, we make the following contributions:
  • We reformulate the problem of fair machine learning as a resource-constrained problem, where we treat the positive decisions as resources to allocate among different groups. We propose an algorithm that enforces fairness under resource limitations.
  • We provide a connection to leveling up, demonstrating that under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
  • We quantify the cost of fairness in this resource-constrained framework. We study the impact of factors inherent to the dataset. We find that the selected positive decision rate has a substantial impact on this cost, an aspect that is not taken into account in previous evaluations.
  • We explore the practical implications of our findings and discuss what strategic actions practitioners might undertake to mitigate trade-offs between groups when fairness is constrained by resource limitations.
2 Background
We consider fairness in classification. Starting with a classifier $c_{w}(\cdot)$ parameterized by weights $w$. Let $\mathrm{H}_{\mathcal{D}}[c_{w}]$be a measure of expected harm of a classifier $c_{w}(\cdot)$ over a particular distribution $\mathcal{D}$ where $y_{x}$ represents the true target label of instance $x$, i.e.,
$\mathrm{H}_{\mathcal{D}}[c_{w}]=\mathrm{E}_{x\in\mathcal{D}}H(y_{x},c_{w}(x))$ (1)
Such measures of a harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$might be $1-\text{precision}$ (for example, when measuring the proportion of people incorrectly stopped by the police); or $1-\text{recall}$ (when measuring the number of cancer cases that go unflagged for follow-up treatment).
In fair classification, we usually want to measure fairness with respect to a protected attribute, such as gender or ethnicity. Using this protected attribute, we can partition the dataset into groups. Group fairness metrics measure (in)equality between these groups with respect to some measure of harm. For example, equal opportunity requires the recall between groups to be equal, while demographic parity enforces an equal selection rate (Verma and Rubin, 2018). Typically, a fair classifier is found by minimizing some global loss $\ell$ (such as accuracy or a continuous proxy such as the logistic loss) while ensuring that the harm is the same per group.
This means that we are searching for a solution to the following problem (where $\mathcal{G}$ is a partitioning of the distribution into groups with respect to a particular protected attribute):
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))\text{ such that }$ (2)
$\mathrm{H}_{g_{1}}[c_{w}]=\mathrm{H}_{g_{2}}[c_{w}]\,\forall g_{1},g_{2}\in\mathcal{G}$
Typical notions of fairness include demographic parity, where the harm corresponds to $1-$ the selection rate, and equal opportunity where it corresponds to the false negative rate.
2.1 Leveling Down
Mittelstadt et al. (2023) observe that methods to enforce fairness often level down; that is, they may enforce fairness by decreasing harm in some groups, but also by increasing harm to other groups (e.g. naively enforcing Equal Opportunity while detecting cancer will often result in a higher false negative rate for some groups i.e. they receiving a lower rate of cancer detection than they would otherwise). This process of enforcing fairness can alter the overall selection rate of the model (Goethals et al., 2024), and is distinct from any leveling down that might be inherently required by the resource constraints discussed earlier. To this end, they suggested replacing Equation LABEL:eq:fair_class with rate constraints, which enforced that the harm in question (e.g. being denied follow-up care) should be below a certain level $k$ for every group:
$\min_{w}\,\mathrm{E}_{x\in\mathcal{D}}\ell(y_{x},c_{w}(x))\text{ such that }\mathrm{H}_{g}[c_{w}]\leq k\,\forall g\in\mathcal{G}$ (3)
This is the same as saying that in the worst case, the harm should be below k, for any group. The challenge now becomes, “How should $k$ be set?” When deciding to deploy a particular model, stakeholders and data scientists often have to agree on acceptable global levels of harm, for example an acceptable recall rate (referred to as sensitivity in the medical literature) for early cancer detection. A similar process can be used to select $k$ instances per group.
This notion of leveling-up, and decreasing the harm in Equation 3 is related to but distinct from minimax-fair machine learning, which minimizes the error for the worst-off group (Martinez et al., 2020). In both minimax fairness and leveling-up the concern is with reducing the harm to the worst-off group and only increasing harms to other groups when necessary, but in minimax fairness the harm is exclusively assumed to be caused by a lack of accuracy or high log-loss (this leads to a formulation of minimize the maximal loss, hence minimax). In comparison, leveling-up is concerned with decreasing other harms such as per-group recall or selection rate, which are distinct from accuracy.  [^3] Both notions are closely aligned to the philosophical principle of the ‘maximin rule’ (Rawls, 2017). According to Rawls, resources should be distributed such that they maximize the benefits for the least advantaged members of society. However, as a fundamental limitation, neither minimax nor leveling-up considers what should be done under constrained resources.
Taking a step back, we can ask the questions: “If $H$ is a harm, why not set $k$ to zero? If particular groups are being harmed by not being selected, why not select everyone, and not bother with machine learning? Similarly, if people are being harmed by a failure to detect cancer, why not schedule everyone for follow-up testing?”
There are multiple possible answers here. One is a matter of personal utility, that harms are generally not one-sided. Failure to repay a loan can lead to bankruptcy for both the lender and customer, resulting in devastating personal consequences. Scheduling unnecessary medical tests is at best alarming, and in the worst case can result in death, depending on how intrusive the follow-up tests are. These types of considerations have been addressed by prior work on the cost of fairness, where personal utility is measured by accuracy, (Menon and Williamson, 2018) and fairness on the ground, which measures the effect of the actual implementation of fairness principles in real-world situations (Bakalar et al., 2021). Constrained resources provide another answer. Ideally, we would prefer to offer scholarships to every student and fast-track treatment for every patient. However, real-world scenarios often lack the necessary resources to achieve this. Understanding how fairness choices are limited and guided by resource constraints is thus crucial to move fair machine learning from a testing to production environment.
Fair classification under limited resources is also related to the domain of fair allocation within welfare economics. This field focuses on ensuring that resources such as time or physical goods are distributed amongst actors in a way that meets certain criteria. This leads to a tension between maximum utility and fairness (Bertsimas et al., 2011, 2012; Donahue and Kleinberg, 2020). Elzayn et al. (2019) describe a resource allocation problem as a problem where there is a limited supply of resources to be distributed across multiple groups with different needs, such as allocating loans or disaster response. We suggest that when the decision-maker has limited resources, all fair classification tasks can be interpreted as resource allocation problems, where positive decisions represent the resources to be distributed.
2.2 Constrained Resources
We formalize the notion of resources, as the number of instances predicted as positive by the machine learning model. This term can be used interchangeably with capacity, or with selection rate, positive decision rate or positive prediction rate when speak about the proportion of positively predicted instances. We define $\mathrm{r}_{\mathcal{D}}[c_{w}]$to be the selection rate (i.e. the expected proportion of positive decisions) of a classifier over a distribution $\mathcal{D}$.
Proposition 1.
Under constrained resources, leveling up results in a solution that enforces equality in harm between groups.
Proof.
We consider the problem of optimizing the distribution of limited resources to minimize the maximum harm experienced by any group, given by:
$\min\max_{g}\mathrm{H}_{g}[c_{w}]\text{ such that }\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$ (4)
where $\mathrm{r}_{\mathcal{D}}[c_{w}]$is a weighted sum of group-specific rates $\mathrm{r}_{g}[c_{w}]$ with all weights positive, thus making $\mathrm{r}_{\mathcal{D}}[c_{w}]$strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$.
We first consider the case where $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing as a function of $\mathrm{r}_{g}[c_{w}]$ (for example, if the harm is $1-\text{recall}$, we expect the harm to be strictly decreasing when more individuals of the group are selected). As $\mathrm{H}_{g}[c_{w}]$ can be written as an invertible function of $\mathrm{r}_{g}[c_{w}]$, $\mathrm{r}_{g}[c_{w}]$ is also strictly decreasing with respect to $\mathrm{H}_{g}[c_{w}]$. Since $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with $\mathrm{r}_{g}[c_{w}]$, and $r_{D}$ is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$, it follows that $\mathrm{H}_{g}[c_{w}]$ is strictly decreasing with respect to $r_{D}$.
We define $W$ as the set of worst-off groups such that:
$W=\arg\max_{g}\mathrm{H}_{g}[c_{w}]$ (5)
The optimum must occur when $\mathrm{r}_{\mathcal{D}}[c_{w}]=r$. If not, one could distribute the remaining resources such that the rates $\mathrm{r}_{g}[c_{w}]$ for the worst-off groups increase without exceeding r, reducing the maximum harm (since $\mathrm{H}_{g}[c_{w}]$ is decreasing with respect to $\mathrm{r}_{g}[c_{w}]$).
Furthermore, the optimum is reached when the harm levels are equal across all groups (so all groups are in $W$). If this were not the case, then there would be at least one group not in $W$, that could sacrifice some of its resources to the worse-off groups. Adjusting the allocation in such a way that the group not in $W$ receives less, while groups at maximal harm receive more, would lead to a reduced overall maximum harm due to the strict decrease of $\mathrm{H}_{g}[c_{w}]$ with $\mathrm{r}_{g}[c_{w}]$. Thus, under these conditions, equality in harm distribution is enforced in the optimal solution, ensuring that the distribution of resources maximally benefits the least advantaged groups within the constraints set by $\mathrm{r}_{\mathcal{D}}[c_{w}]\leq r$.
The proof also holds when the harm considered is strictly increasing with respect to $\mathrm{r}_{g}[c_{w}]$ (e.g., a harm corresponding to $1-\text{precision}$). Here, the optimal solution is trivial with an overall selection rate of 0, and equality will be preserved. However, in this case, it might be more realistic to flip the sign and require the overall selection rate to be at least some r. We can simply replace the classifier $c_{w}$ with its negation, and now find that the harm is strictly decreasing with respect to the selection rate of the new classifier. Revisiting the previous proof, we find that equality holds.
Hence, we find that equality must hold in all four cases (strictly increasing or strictly decreasing harm, and a global selection rate that is either constrained above or below some value). ∎
2.3 Cost of Fairness
Before we proceed further, we should note that assigning a cost to actions is inherently a political action, and that often these costs reflect the beliefs of data scientists and other professionals as much as they do the raw data. This is particularly apparent in the case of personal loans.
While banks have good knowledge of the repayment habits of people they have lent money to in the past, they have much less knowledge of how people they did not lend money to might repay. In no small part, this knowledge asymmetry has allowed for the persistence of institutional redlining whereby loans are routinely denied to people in black or Hispanic dominant neighbourhoods when they would be granted to individuals in similar circumstances that live in white dominant neighbourhoods  (Markup, 2021). Similarly, while it is tempting to assume that medical data is reliable, and for example, that we are forecasting the result of a future biopsy that will infallibly determine if cancer is present or not, for many problems the information is less clear-cut, and ground-truth may come from two experts attempting to diagnose from limited post-mortem records, with a third expert adjudicating in the case of disagreement Lee et al. (2023).
Indeed, one justification for constraints such as demographic parity is that ground-truth data is often gender- or ethnically-biased, and in some circumstances we can a priori expect uniform rates across these populations (in line with the “We’re All Equal” world-view of Friedler et al. (2021), which asserts that there are no innate differences between groups). As such, any estimate of cost comes with the usual caveat of assuming the ground-truth is correct. But even as simplified approximations, these costs remain useful for understanding the potential trade-offs in enforcing fairness.
With this in mind, we consider the following question: What is the global change in harm from an optimal classifier when we minimise the harm of the worst-off group?
Under constrained resources, this leads to the same results as enforcing fairness, i.e. equality of harms between groups. We can measure this by the difference between the harm of the optimal classifier $c^{o}_{w}(.)$ and the classifier that satisfies fairness $c^{f}_{w}(.)$.
$\mathrm{H}_{\mathcal{D}}[c^{f}_{w}]-\mathrm{H}_{\mathcal{D}}[c^{o}_{w}]$
While other studies have previously analyzed the cost of fairness (Friedler et al., 2019; Haas, 2019; von Zahn et al., 2021), they did not consider the implicit trade-off that comes from constrained resources in decision-making, but instead measured how classifiers deteriorate with fairness.
We might also be interested in performance metrics outside the fairness metric enforced. To quantify the loss in precision, we ask: What is the change in precision if we minimize the harm of the worst-off group? This represents a common scenario. In a medical context it corresponds to the question: What proportion of healthy patients instead of sick will we see if we increase test sensitivity for the worst off groups? In the results, we measure this cost and its contributory factors. This leads to a follow-up question: What is the increase or decrease in selection rate needed to preserve the current rate of global harm, if we minimise the harm of the worst off-group? The last question represents a common political solution to this problem. When particular groups are disadvantaged by the status quo, it is often easier to increases the resources available, and target these resources at the disadvantaged groups, rather than requiring currently advantaged groups to accept less access to resources. The number of available resources can be an alterable parameter that should explicitly be taken into account.
3 Resource-constrained Fairness and its Cost
While standard fairness methods do not enforce fairness as the selection rate varies (see Figure 2), there is a straightforward algorithm for enforcing fairness at any selection rate, based on the harms based analysis of section 2.2. Given an existing classifier $c_{w}$, we say that a classifier makes a positive decision regarding datapoint $x$ if $c_{w}(x)>t$, where t is some threshold. We consider fairness as in Equation LABEL:eq:fair_class where some notion of harm $\mathrm{H}_{\mathcal{D}}[c_{w}]$is being equalized across groups.
Now, we simply vary a per group threshold $t_{g}$, while computing the group harm $\mathrm{H}_{g}[c_{w}]$, and the corresponding selection rate $\mathrm{r}_{g}[c_{w}]$. We consider some $h$, and select the set of thresholds so that the harm per group lies just under this $h$. This is (approximately) fair with respect to harm $H$, and has a global selection rate
$r=\frac{\sum_{g}|g|\,\mathrm{r}_{g}[c_{w}>t_{g}]}{\sum_{g}|g|}$ (6)
where $|g|$ is the number of datapoints in group $g$. By sweeping over possible values $h$, and examining the corresponding global selection rate, we can find a fair solution that lies just under a target selection rate. The corresponding thresholds can be used to enforce fairness on previously unseen test data.
The resource allocation process is particularly straightforward for demographic parity. In this case the harm corresponds to 1- the selection rate, and the process simplifies to just selecting the top $r$ from each group.  [^4] Note that in particular case of demographic parity ideal thresholds can be calculated on the test set, but for most other measures the thresholds need to be calculated on a separate validation set, as they require access to the target label in order to compute the harms. Owing to sampling error, thresholds selected on the validation data do not perfectly correspond to selection rates on unseen data, and to ensure that the number of instances predicted as positive exactly matches the available resource level, we do not directly use the thresholds. Instead, we calculate the proportion of resources to allocate to each group for each resource level $R$ on the validation set, and then enforce this on the test set.
Our strategy belongs to the class of post-hoc bias mitigation strategies, that attempt to make the output of machine learning model fair after the model has been trained (Mehrabi et al., 2021). We conduct this analysis across a complete range of selection rates $r$ (from 1% to 100 %) to represent different resource levels.
Given this approach, we can now quantify the cost of fairness, defined as the difference in performance (measured as either precision or recall) between the default and fair labeling approaches, for a fixed number of selected instances (R). We vary the available resources $R$ (and consequently the selection rate $r$) and investigate how this influences the cost of fairness.
3.1 Datasets and Classifiers
We use a standard train-test split, where the model is trained on the training set and the predictive performance and cost of fairness is evaluated on a separate test set.
We use several real-world datasets common in tabular fair machine learning (Le Quy et al., 2022). We also extend our findings to two datasets (CelebA Liu et al. (2015) and Fitzpatrick17k Groh et al. (2021)) from computer vision. The Adult Income dataset is taken from from the 1994 census data, and contains a prediction task identifying if an individual’s annual income surpasses \$50,000. The Compas dataset gathers demographic details and criminal records of defendants from Broward County to forecast the likelihood of reoffending within two years. The Dutch Census dataset from 2001 captures aggregated demographic data in the Netherlands, utilized to determine if an individual’s occupation falls into a high-level (prestigious) or low-level category. The Law Admission dataset contains data from a 1991 survey by the Law School Admission Council (LSAC) across 163 U.S. law schools, aimed at predicting a student’s success on the bar exam. The CelebA dataset comprises over 200,000 celebrity images, each annotated with 40 attribute labels ranging from hair color to emotions, along with 5 landmark locations. We use the attribute ‘wearing earrings’ as target label, as it is highly skewed towards the non Male class Wang et al. (2020) and thus has a significant base rate disparity. The Fitzpatrick17k dataset consists of approximately 17,000 dermatologist-curated skin lesion images, categorized across the Fitzpatrick skin type scale. We prepossess the data and binarize the labels following the approach of Zong et al. (2022). We specify the protected and target attributes in Table 1. We also report the base rate disparity for each dataset, which is defined as the actual difference in the proportion of positive outcomes between groups in a dataset. If we consider a binary sensitive attribute $S$, with $ns$ representing the protected group and $s$ the advantaged group it can be mathematically expressed as:
$P(Y=1\mid S=ns)-P(Y=1\mid S=s)$
Table 1: Used datasets. The values between brackets represent the percentage of the dataset with that value.
[…]
Classifiers
For the tabular datasets, we use eXtreme Gradient Boosting (XGBoost) Chen and Guestrin (2016). We optimize the number of boosting rounds through 5-fold cross-validation on the training set with early stopping after 10 rounds if no improvement is seen. The optimized model is trained on the entire training set and evaluates the probability of positive class outcomes. For the image datasets, we employ a Resnet-50 (CelebA) and Resnet-18 (Fitzpatrick-17k) backbone He et al. (2016) pretrained on ImageNet Deng et al. (2009) for feature extraction.
3.2 Metrics
3.2.1 Performance metrics
Each classifier returns a prediction score $S$, where higher $S$ values indicate a greater likelihood of $Y=1$. This score is converted to a binary prediction using a threshold: $\hat{Y}=\mathds{1}\{S>t\}$. Evaluating classifier performance with the prediction scores, unaffected by the threshold choice, is best measured by the Area Under the ROC Curve (AUC). The AUC score is given by:
$P(S(x_{i})>S(x_{z})\mid y_{i}=1,y_{z}=0)$
This formula measures the probability that the classifier ranks a randomly chosen positive instance higher than a randomly chosen negative instance.
We evaluate the performance of the default allocation and the fair allocation over a range of selection rates (where the number of selected instances is the same for both allocations). When the selection rate of a machine learning model is fixed, it implies that a set proportion of the instances will be chosen based on the highest scores predicted by the model, regardless of their actual scores. A useful performance metric to measure is the proportion of relevant instances among the top $R$ instances selected by the model (where $R$ denotes the resource level). This means the number of selected instances that actually have a positive target label.
Precision at R=Number of actual positives in the top $\text{Precision at $R$}=\frac{\text{Number of actual positives in the top $R$}}{R}$
While precision at the top $R$ relates to the efficient use of limited resources, knowing this value also determines the values of both recall and accuracy at $R$ for a specific dataset and model and a fixed value of $R$ (Rodolfa et al., 2021). We measure the cost of fairness as the loss in precision when using the fair allocation instead of the optimal allocation. For the sake of completeness, we also report the cost as loss in recall [^5] The proportion of actual positives in the top $R$ over the total number of actual positives in Table 2 and Figure 5.
3.2.2 Fairness Metrics
We focus on two widely recognized fairness measures used to assess disparities between groups.
Demographic parity (also known as statistical parity) requires that the rate of positive decisions is roughly equal for both the protected group and the privileged group:
$P(\hat{Y}=1|S=s)\approx P(\hat{Y}=1|S=ns)$
Equal opportunity requires the true positive rate to be approximately the same across groups (Hardt et al., 2016), which enforces equal recall:
$P(\hat{Y}=1|S=s,Y=1)\approx P(\hat{Y}=1|S=ns,Y=1)$
4 What influences the Cost of Fairness?
Table 2: Model performance (entire population, privileged group and protected group) and the average cost of fairness (the loss in precision and recall) when enforcing DP and EO. The average is calculated over all the selection rates from $1\%$ to $100\%$.
Dataset Adult Compas Dutch Law CelebA Fitzpatrick17K
AUC 0.899 0.811 0.917 0.871 0.957 0.826
$AUC_{priv}$ 0.884 0.797 0.884 0.852 0.931 0.829
$AUC_{prot}$ 0.887 0.793 0.914 0.842 0.969 0.814
Avg. loss in precision (DP) 0.024 0.025 0.028 0.006 0.069 0.001
Avg. loss in precision (EO) 0.010 0.013 0.007 0.004 0.009 0.000
Avg. loss in recall (DP) 0.024 0.016 0.028 0.004 0.062 0.001
Avg. loss in recall (EO) 0.012 0.007 0.008 0.003 0.013 0.001
In Table 2, we analyze the AUC performance of our machine learning model for different groups: the entire population, and separately for protected and privileged groups. Additionally, we report the average cost of enforcing fairness as the loss in both precision and recall between the default allocation and the fair allocation, averaged over all selection rates. We see that the average cost of enforcing fairness is the lowest for the Fitzpatrick17K dataset, which also has a very low base rate disparity. The average cost is also very low for the Law dataset, despite its significant base rate disparity. A possible reason for this could be the relatively small size of the protected group in this dataset, which results in less reallocation of resources. The average cost of enforcing fairness across other datasets is more similar, however it is worth noting that the cost associated with enforcing both fairness metrics on CelebA is higher than for the other datasets. It is hard to attribute the difference in costs between the datasets to a single factor, as many of the parameters will be different. This is why we perform a separate analysis on the Adult Income dataset.
Here, we systematically vary key parameters—one at a time—to observe their effects on the cost of fairness. Specifically, we explore modifications to the base rate disparity, the size of the protected group, the noisiness of the whole dataset and the noisiness of the protected group. The impacts of these changes are visualized in Figure 2, with the original dataset’s results represented by a black line in each figure.
(a) Base rate disparity (DP)
(b) Base rate disparity (EO)
(c) Noise (DP)
(d) Noise (EO)
(e) Subgroup noise (DP)
(f) Subgroup noise (EO)
(g) Subgroup size (DP)
(h) Subgroup size (EO)
Figure 2: Impact of parameters on the cost of fairness, measured as the loss in precision (Adult Income dataset). We see that reducing the base rate disparity leads to a lower cost of fairness. Similarly, introducing more noise across all groups, also lowers the cost of fairness, but introducing more noise exclusively in the protected group, increases the cost of fairness. Finally, reducing the size of the protected group, results in a decrease in the cost of fairness as well.
(a) Adult
(b) Compas
(c) Dutch
(d) Law
(e) CelebA - Earrings
(f) Fitzpatrick17k
Figure 3: Cost of fairness (loss in precision) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). The cost of fairness for both metrics heavily depends on the available resource level and thus the used selection rate.
Base rate disparity
First, we alter the base rate disparity. We see in Figures 2(a) and 2(b), that a reduction in the base rate disparity leads to a lower cost of fairness. This trend is more pronounced when enforcing DP compared to EO, but it is evident under both fairness metrics. Menon and Williamson (2018) also find that the trade-off between accuracy and fairness is determined by the strength of the correlation between the sensitive attribute and the target.
Noise
We alter the performance of the model by introducing random noise (with varying degrees from 0 to 0.5) to the feature values of the whole dataset. The base rates of both groups remain the same, but the model struggles more with correctly classifying individuals, resulting in a lower AUC score. We see that as model performance decreases, the average cost of fairness also goes down, both for DP and EO (Figures 2(c)-2(d)). This global decrease shrinks the gap between the performances of each group, reducing the overall cost of fairness. This phenomenon might also explain why the average cost of fairness is so high for the CelebA dataset in Table 2. The model is very good in distinguishing negatives and positives from each other (as measured by the AUC), so the average cost will be high. [^6] Following this reasoning, we see that if we would fit a perfect model (so the model predicts the target label perfectly for all instances) to a dataset with some level of base rate disparity, then the maximum cost would be reached at a point between the base rates and be exactly equal to the level of base rate disparity in that dataset. In the case of CelebA, the model is not perfect but the maximum cost is also relatively close to the level of the base rate disparity, and is reached at a point between the base rates.
Subgroup noise
We can also add noise only to the members of the protected group. This will lower the performance of the model for individuals of the protected group, but not for individuals of the privileged group. As shown in Figures 2(e) and 2(f), this modification increases the cost of fairness for both DP and EO, with a larger effect for EO. Chen et al. (2018) also find that when there is a difference in noise level, and available covariates are not equally predictive of the outcome in both groups, fairness cannot be satisfied without sacrificing accuracy. Dutta et al. (2020) confirm that if there is not enough separability information for one group compared to the other, being fair will reduce the accuracy.
Size of the protected group
Lastly, the size of the protected group substantially influences the cost of fairness. When the protected group is smaller, fewer adjustments are necessary to achieve fairness, thereby reducing the cost. This relationship is clearly illustrated in Figures 2(g) and 2(h), where reducing the size of the protected group, by randomly excluding part of this group, consistently lowers the cost of fairness. This is an expected finding, however, we have not seen it discussed previously.
4.1 The impact of selection rate
Figures 3(a)-3(f) demonstrate the loss in precision for various selection rates when using a fair allocation compared to the default allocation of the machine learning model. We see that this cost varies substantially depending on the available resources. For the Adult (Figure 3(a)), Compas (Figure 3(b)), CelebA (Figure 3(e)) and Fitzpatrick 17K (Figure 3(f)) datasets, the cost is highest for low selection rates, while for the Dutch (Figure 3(c)) dataset, the cost is the highest for medium selection rates, and for the Law (Figure 3(d)) dataset, the cost is the highest for high selection rates. The influence of resource level on the cost of fairness has not been discussed before, yet it is a critical factor. Depending on the resource level, the cost can either be negligible or very high. The loss in recall can be seen in Figure 5. We see that the patterns are in line with the loss in precision. We can also partially agree with Rodolfa et al. (2021) who posited that for very constrained top-$k$ settings, so when the selection rate is very low, the cost of fairness can be negligible. However, we only find this to be true when the base rate is a lot higher and the performance of the model is good enough. When the base rate is low, the cost of fairness will usually be high for low-resource settings.
(a) Selection rate = 10 %
(b) Selection rate = 25 %
(c) Selection rate = 50 %
(d) Selection rate = 75 %
Figure 4: Allocation results for different resource levels (Adult Income dataset). Left point on the x-axis represents the resources being maximally allocated to the privileged group, while the right point represents the resources being maximally allocated to the protected group.
We can get more insights into how these trade-offs behave for each resource level by analyzing the different scenarios. For the Adult dataset, this is shown in Figures 4(a)-4(d). These figures demonstrate how different allocations lead to different levels of precision for one fixed resource level. The two extreme points of the curve represent all the resources being awarded to either the privileged group (red dot on the left) or the protected group (purple dot on the right).  [^7] However, also for these extreme points, when the total level of resources is higher than the size of one of the groups, some of the resources will still be awarded to the other group (as can for example be seen in Figure 4(c)). We see that the optimum is not reached by awarding all the resources to one of the groups, but somewhere in the middle (darkblue dot). We note that the precision of the unconstrained (‘unfair’) model, will be very close to the optimal allocation, and that the allocations required by the fairness metrics (both DP and EO) will lead to a lower precision for every resource level (Figures 4(a)-4(d)). However, this difference is a lot larger for low resource levels (Figures 4(a)-4(b)). In Figure 4(c), the default allocation of the machine learning model is very unfair, but using the fair allocation (both DP and EO) results in only a little change in precision. This demonstrates that a very unfair initial model allocation does not guarantee a high cost of fairness. In Figure 4(d), the default allocation of the machine learning model is already approximately fair. Hence, the cost of fairness is also low.
4.2 The Impact of the Fairness Metric
We can see in Table 2, and in Figures 34 and 5 that the cost of fairness is usually lower when enforcing equal opportunity than when enforcing demographic parity. This is because the number of instances that needs to be ‘swapped’ is typically lower when enforcing equal opportunity. Liu et al. (2018) also find that the selection rate enforced by equal opportunity is likely to be much closer to the optimal than the selection rate enforced by demographic parity, while Hardt et al. (2016) show that enforcing equal opportunity leads to less loss in profits than enforcing demographic parity.
The setting where EO would results in a higher number of swapped instances, is when an equal true positive rate results in stricter requirements than an equal positive rate. A case where this could happen is when one of the groups has significantly lower data quality. DP suffers less under this scenario, as here we do not care if the predictions are accurate or not. However, to enforce EO, the algorithm must do equally bad on all groups, leading to a worse overall performance. Indeed, as shown in Figures 2(e) and 2(f), the cost of enforcing EO increases significantly more rapidly compared to the cost of enforcing DP when one subgroup becomes noisier.
4.3 Bounding the Cost of Fairness
We now formalize these costs and consider the cost of the change induced by enforcing fairness at a particular selection rate.
As we are modeling the cost of inducing fairness under rate constraints, we consider the changes in the labeling induced by swaps, where the label assigned to one instance changes from positive to negative, and the label assigned to another instance changes from negative to positive. A sequence of swaps must always preserve the selection rate, and any pair of labellings with the same selection rate, can be transformed from one to another by a sequence of swaps.
We write $p$ for the proportion of instances in the dataset swapped from positive to negative (or equivalently the proportion swapped from negative to positive). The overall cost of inducing fairness can be written as:
$\text{cost}=p\times\text{average cost of a swap}$ (7)
Note that $p$ is somewhere between 0 and 0.5, and for standard linear measures (e.g. accuracy, recall, sensitivity, and False Positive and Negative Rates) the second term lies somewhere between $-c^{-1}$, and $c^{-1}$. Here,
$c=\frac{\text{\#instances used in the measure}}{\text{\#instances in dataset}}$ (8)
i.e. $c=1$ for accuracy, and, writing $b$ for the base rate, $c=b$ for Recall and False Positive Rate, and $c=(1-b)$ for Specificity and False Negative Rate. As such, $c^{-1}p$, and half the average cost of a swap are upper bounds on the cost of enforcing fairness.
To state this less formally, the cost of fairness, in terms of change in accuracy, is at most the proportion of swaps ($p$); while the cost in terms of change of recall is at most the proportion of swaps divided by the base rate.
We can create successive bounds based on this identity.
Given a fixed global selection rate $r$, $p$ is bounded above by $r$ (all positive instances are swapped); and $(1-r)$ (all negative instances are swapped). Therefore, the cost of fairness is bounded above by $rc^{-1}$, and $(1-r)c^{-1}$. As such, if $r$ or $(1-r)$ is much smaller than $c$, there is little cost in enforcing fairness.
This matches our empirical finding that the cost is bell-shaped with respect to the global selection-rate. The highest costs occur when the selection rates are closest to the selection rates of an unconstrained classifier. This means that many of the existing works on the cost of fairness (Menon and Williamson, 2018; Friedler et al., 2019; Hort et al., 2023) overestimate the cost of fairness for resource-constrained scenarios. For scenarios where the selection rate substantially exceeds the base-rate, this is because for any informative classifier, the majority of candidates that might be positive would be selected early on, regardless of which group they belong to, and the remaining pool of candidates are likely to have negative labels regardless of which group they belong to.
Similarly, given $g$, the proportion of the dataset occupied by the smallest group, we know $p<g$, and the cost of inducing fairness is bounded above by $gc^{-1}$. [^8] Relabelling the smallest group is not always sufficient to enforce fairness, but it holds for the fairness metrics of demographic parity and equal opportunity that we study in this paper. Again, if $g$ is much smaller than $c$, there is little cost in inducing fairness. This result is supported empirically by our experiments (see Figures 2(g) and 2(h)).
Finally, looking at the second term (the average cost of relabelling), we find this second factor explains much of what we see. For example, increasing global noise decreases performance for both groups, at broadly similar proportions. This overall decrease shrinks the gap between the per-group performance, reducing the average cost of relabelling points. Similarly, increasing classification noise for the group with greatest classification error, or increasing the difference in base rates, increases the performance gap between groups and consequentially the cost of relabelling.
Bounding loss in precision
These bounds are much weaker for non-linear measures such as precision, if they exist. For precision, assuming a fixed global selection rate $r$, we have $c=r$. Consequentially, the low selection-rate bound $rc^{-1}=1$ is particularly meaningless. However, other bounds remain informative when there is a higher selection rate – particularly, the bounds $(1-r)c^{-1}$ and $gc^{-1}$, indicate that we should see similar patterns in the costs with respect to precision for higher selection rates. Empirically, we find this — precision exhibits similar trends to the linear measures of recall and accuracy. Reflecting the much weaker bound for low selection rates, it sometimes exhibits strong changes at low resource levels, which are not present for accuracy and recall. This can be seen in Figure 3(d).
4.4 Varying Global Selection Rates
Strategic adjustments in selection rate may be necessary depending on the outcomes of a cost-benefit analysis. Until now, we assumed the resource level to be fixed, but practitioners may find it feasible to increase or decrease this level, particularly if the resource levels were initially determined under an unfair model allocation. For instance, in healthcare, additional investments could be directed towards expanding screening facilities to counterbalance the decrease in true positives, effectively maintaining the detection rates prior to implementing fairness measures. Similarly, in educational settings such as student admission, institutions might adjust the size of admitted cohorts, either decreasing it to preserve the average quality, or increasing it to have the same number of graduates. Donahue and Kleinberg (2020) also discuss how increasing the level of available resources is a critical goal where advocacy and political action can play a key role.
Following section 4.3, it is reasonable to ask if bounds exist for levelling up, and if there are bounded increases in selection rate for a disadvantaged group guaranteed to remove unfairness (while keeping the selection rate of the advantaged group as-is). In brief, the answer is yes for demographic parity; for equal opportunity you may need to relabel all the data in a disadvantaged group; and levelling up is not always possible for equal precision. These results are so different from section 4.3 because the analysis of the previous section was advantaged by the stability of measures – we altered per group selection rates and hoped that global measures remained stable. In this new case, we are trying to alter per-group measures and the stability of measures is now a disadvantage.
For demographic parity, there are no stability concerns. Let $g$ be the proportion of the dataset corresponding to the disadvantaged group, and $r$ the global classifier selection rate. Then an upper bound for the selection rate excluding members of the disadvantaged group is $r/(1-g)$, and at most we require an additional $gr/(1-g)$ proportion of points to be positively labelled to obtain parity.
For equal opportunity, only instances assigned a positive ground-truth label alter the recall, and it is possible that a badly-trained classifier selects all negatively labeled instances first, and only then positively labelled points. If we consider a classifier like this, and a group with only one positively labelled point (in the ground truth), all points within the group must be selected to get any non-zero recall rate for that group. As such, $g$ is a bound.
For equal precision, we consider the same example, and note that while you must select all datapoints to obtain non-zero precision, the precision will be the base rate of the group, and it may be insufficient for equality with respect to the other group.
Another strategic approach could involve refining the precision for the protected groups by collecting more data or enhancing data quality, albeit at an initial cost. This investment might reduce the overall fairness cost, yielding a more profitable model over time. However, high quality data may simply not be available for all groups or infeasible to collect for a variety of legal, political, and practical reasons, meaning it cannot be assumed that collecting more or “better" data will necessarily solve inequalities in resource distribution (Perez, 2019; Pot et al., 2019; Wachter et al., 2020).
5 Conclusion
This work addresses the practical concern of enforcing fairness with a limited budget or, equivalently, as the selection rate/threshold varies. We demonstrated that standard methods for fairness do not enforce fairness for a varying threshold (Figure  2). We unified leveling up (Mittelstadt et al., 2023) and fairness under budgetary constraints, and have shown that the solutions for each approach coincide under constrained resources (Section 2.2). We have also shown how simple methods for setting per group thresholds can enforce fairness under these rate constraints (Section 3). Using this, we empirically measured the cost of fairness under rate constraints – in terms of decreased performance, or as an increase in global harm (Section 4); and derived theoretical bounds for these costs (Section 4.3). While the bounds are not intended to be tight, they are informative, and align with our empiric findings.
This allows us to investigate the actual cost of fairness, rather than changes in performance metrics that are heavily influenced by a change in the overall selection rate (Goethals et al., 2024). Compared to other fairness analyses, our analysis more closely aligns with real-world challenges, where resources are often fixed, and allows organizations to have reasonable expectations about what costs to expect and why. Furthermore, we demonstrated that the decision-making context matters, particularly the behavior of the classifier; the used fairness metric; and the level of available resources. This highlights the importance of evaluating the cost of fairness within the specific domain of use.
Acknowledgments
This work has been supported through research funding provided by the Wellcome Trust (Grant number 223765/Z/21/Z), Sloan Foundation (Grant number G-2021-16779), Department of Health and Social Care, EPSRC (Grant number EP/Y019393/1), and Luminate Group. Their funding supports the Trustworthiness Auditing for AI project and Governance of Emerging Technologies research programme at the Oxford Internet Institute, University of Oxford. Sofie Goethals was funded by Flemish Research Foundation (Grant number 11N7723N), and by a travel grant (Grant number V413024N) awarded by Flemish Research Foundation to visit the University of Oxford.
Appendix
(a) Adult
(b) Compas
(c) Dutch
(d) Law
(e) CelebA - Earrings
(f) Fitzpatrick17k
Figure 5: Cost of fairness (loss in recall) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). We see that the cost of fairness for both metrics heavily depends on the available resource level and thus the used selection rate.